OpenStudy (anonymous):

HELP- solve the equation: 48n^2 + 30n = 0

OpenStudy (anonymous):

factor out common things. They both can factor out 6n which leaves you with 8n + 5 = 0 then move 5 over and divide by 8

OpenStudy (anonymous):

n=-5/8 and x= 0

OpenStudy (anonymous):

n=0*

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

hook me up with a medal :p